An object has a kinetic energy of 254 J and a momentum of magnitude 23.7 kg·m/s. Find the speed and the mass of the object.

speed m/s
mass kg
Answer #1

KE = P^2/2m ( where P is momentum and m is mass of partice)

=> m = P^2/(KE) = 23.7^2/(254) = 2.211 kg

speed = momentum/mass = 23.7/2.211 = 10.71 m/s